# DM-Codec: Distilling Multimodal Representations for Speech Tokenization

### v1

@@ 4.1 Comparison of Speech Tokenization models
Table 2: Significance Analysis of DM-Codec (D) compared to baselines EnCodec (E), SpeechTokenizer (S), and FACodec (F). Results reveal DM-Codec consistently achieves significantly better scores in key metrics across all individual samples.   indicates that DM-Codec is significantly better, a   denotes dominance, and a   means no significant improvement over the baseline. Avg and Std mean the average and standard deviation of each score.
@@
DM-Codec
Avg Std E S F Avg Std E S F Avg Std E S F Avg Std E S F
0.053 0.113 0.082 0.157 3.258 0.184 0.937 0.019
EnCodec
Avg Std D S F Avg Std D S F Avg Std D S F Avg Std D S F
0.061 0.131 0.090 0.158 3.078 0.201 0.920 0.017
SpeechTokenizer
Avg Std E D F Avg Std E D F Avg Std E D F Avg Std E D F
0.060 0.139 0.089 0.166 3.087 0.190 0.923 0.021
FACodec
Avg Std E S D Avg Std E S D Avg Std E S D Avg Std E S D
0.057 0.123 0.086 0.163 3.129 0.250 0.949 0.923
4.2 Significance Analysis of Speech Tokenizer performance
[…]
@@ 4.2 Significance Analysis of Speech Tokenizer performance
Results and Discussion: The results in Table 2 show that DM-Codec significantly outperforms the baselines in WER, WIL, ViSQOL, and STOI scores. The improved average values (0.053 WER, 0.082 WIL, 3.258 ViSQOL, 0.937 STOI) and consistent standard deviations (0.113 WER, 0.157 WIL, 0.193 ViSQOL, 0.019 STOI) further demonstrate the statistical significance. Notably, DM-Codec’s performance in WER and WIL underscores the importance of contextual representation distillation for enhanced speech reconstruction. Additionally, its dominance in ViSQOL and STOI, especially over EnCodec, highlights the benefits of combining LM and SM distillation for retaining semantic-acoustic fidelity. While DM-Codec does not achieve significant dominance over FACodec in terms of STOI, it significantly outperforms the baselines across all other metrics. Among the baseline, however, FACodec achieves improved results over EnCodec and SpeechTokenizer, whereas SpeechTokenizer surpasses EnCodec in performance.
4.3 Ablation Studies
We conducted a thorough analysis of DM-Codec’s performance and the impact of each methodological choice in LM-guided and combined LM and SM-guided distillation. Unless otherwise stated, we use distillation for both LM and SM from the first Residual Vector Quantizer layer (RVQ-1) for comparison consistency and simplicity.
4.3.1 Ablation Study: Impact of Combined Semantic Distillation
Table 3: Effects of weights on combined representation distillation: Higher LM weight enhances content preservation, leading to lower WER. $\lambda_{SM}$ is the SM weight, $\lambda_{LM}$ is the LM weight.
@@ 4.3.1 Ablation Study: Impact of Combined Semantic Distillation
$\mathcal{L}_{LS}=\frac{1}{2}\left(\lambda_{SM}\cdot\mathcal{L}_{SM}+\lambda_{LM}\cdot\mathcal{L}_{L}\right)$ (11)
Results and Discussion: The experimental results are presented in Table 3, showing the speech reconstruction results with WER scores for different weighted combinations. From the values, we notice a trend showing that incorporating LM representations significantly improves WER, especially when LM distillation is dominant. The lowest WER score of 4.07 occurs with a weight of $\lambda_{LM}=0.8$ for LM, while $\lambda_{SM}=0.2$ for SM, highlighting the strong influence of LM distillation on capturing contextual information. A balanced weighting of $\lambda_{SM}=0.5$ and $\lambda_{LM}=0.5$ produces a WER of 4.18, confirming that distillation from both LM and SM is beneficial. However, as the weighting shifts more in favor of SM ($\lambda_{SM}>0.7$), WER deteriorates, reaching 4.83 when relying entirely on SM. This underscores that over-reliance on SM distillation compromises contextual accuracy in favor of raw speech features. Thus, an LM-dominant approach yields optimal results, while using SM alone is less effective in preserving content.
4.3.2 Ablation Study: Impact of Distillation on Different RVQ Layers
@@
Table 4: Analysis of different RVQ layers effect on speech reconstruction. LM-guided distillation on RVQ-1 layer ensures greater content preservation, while SM-guided distillation on RVQ-1:8 layer is more effective at preserving semantic representation. LM-layer and SM-layer indicate the RVQ layer used for respective distillation. $\clubsuit$ indicates LM-guided Distillation. $\spadesuit$ indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec $\clubsuit$ RVQ-1 - 4.36 7.06 3.18 0.935
DM-Codec $\clubsuit$ RVQ-1:8 - 4.23 6.94 3.12 0.929
DM-Codec $\clubsuit$ RVQ-8 - 4.44 7.22 3.28 0.935
DM-Codec $\spadesuit$ RVQ-1 RVQ-1 4.18 6.84 3.13 0.933
DM-Codec $\spadesuit$ RVQ-1:8 RVQ-1 4.59 7.34 3.21 0.937
DM-Codec $\spadesuit$ RVQ-8 RVQ-1 4.49 7.24 3.30 0.938
DM-Codec $\spadesuit$ RVQ-1 RVQ-1:8 4.05 6.61 3.26 0.937
DM-Codec $\spadesuit$ RVQ-1 RVQ-8 4.39 7.08 3.33 0.939
4.3.3 Ablation Study: Impact of Different Models on Distillation
@@
Table 5: Analysis of representation distillation from different models. BERT can be effectively combined with HuBERT or wav2vec 2.0, however, ELECTRA in LM-guided distillation outperforms BERT. $\clubsuit$ indicates LM-guided Distillation. $\spadesuit$ indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec $\clubsuit$ BERT - 4.36 7.06 3.18 0.935
DM-Codec $\clubsuit$ ELECTRA - 4.12 6.63 3.10 0.936
DM-Codec $\spadesuit$ BERT HuBERT 4.18 6.84 3.13 0.933
DM-Codec $\spadesuit$ BERT wav2vec 2.0 4.13 6.77 3.15 0.942
DM-Codec $\spadesuit$ ELECTRA wav2vec 2.0 4.70 7.51 3.14 0.933
DM-Codec $\spadesuit$ ELECTRA HuBERT 4.67 7.58 2.94 0.932
4.3.4 Ablation Study: Impact of Different Distillation Layer(s)
@@
Table 6: Analysis of different distillation layers representation on speech reconstruction. Average layer provides more comprehensive representations. $\clubsuit$ indicates LM-guided Distillation. $\spadesuit$ indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec $\clubsuit$ Average 4.36 7.06 3.18 0.935
DM-Codec $\clubsuit$ Last 4.62 7.56 2.95 0.926
DM-Codec $\clubsuit$ $9^{\text{th}}$ 4.75 7.80 2.88 0.925
DM-Codec $\spadesuit$ Average 4.18 6.84 3.13 0.933
DM-Codec $\spadesuit$ Last 4.68 7.55 3.03 0.933
DM-Codec $\spadesuit$ $9^{\text{th}}$ 4.52 7.43 3.00 0.933
5 Related Work
Tokenization Techniques in Speech. Tokenization in speech processing can be broadly categorized into two main approaches: (i) speech encoder-based and (ii) language-based. In the speech encoder-based tokenization approach, a pre-trained speech encoder provides audio representations. These representations are then used to guide the training model, either through an alignment network (Messica & Adi, 2024) or by optimizing specific losses (Zhang et al., 2024a; Liu et al., 2024). The language-based tokenization approach involves processing audio through a speech encoder to obtain discrete representations, or using the corresponding text to feed into a language model. This has been explored by (Hassid et al., 2024), (Wang et al., 2024),(Zhang et al., 2023), and (Zhang et al., 2024b). Recently, LAST (Turetzky & Adi, 2024), explored a language model to tokenize speech toward improved sequential modeling, using the LLM to perform the next token prediction of quantized vectors. However, these approaches significantly differ from our language representations distillation method and do not focus on combining multimodal representations.
Discrete Speech Representation. There are two well-known methods for discrete speech representation: semantic tokens and acoustic tokens. Semantic tokens are derived through self-supervised learning (SSL) techniques for speech (Baevski et al., 2019; Hsu et al., 2021; Chung et al., 2021) and capture abstract, high-level features that relate to general, symbolic aspects of speech, while omitting details related to speaker identity and acoustic characteristics. In contrast, acoustic tokens are obtained using neural audio codecs (Zeghidour et al., 2021; Défossez et al., 2022; Yang et al., 2023) and focus on delivering precise reconstructions of acoustic features. However, recent models (Turetzky & Adi, 2024; Liu et al., 2024; Shi et al., 2024) have shown that speech models based on self-supervised learning (SSL) are effective at extracting acoustic representations where LMs be employed to refine these models further, enhancing their ability to extract more nuanced semantic representations.
Textual Language Models in Speech. Research on speech models, including works by (Nguyen et al., 2023), (Borsos et al., 2023), and (Kharitonov et al., 2022), has focused on utilizing raw audio to extract prosodic features, identify speaker characteristics, and generate audio without depending on textual features or supervision from textual LMs. In contrast, many newer methods have started using audio encoders to transform audio signals into discrete tokens, which can be processed by textual LMs. TWIST method introduced by (Hassid et al., 2024) initializes the weights of the SpeechLM using a pre-trained text LM, showing that this combination significantly improves performance. Similarly, the SELM model developed by (Wang et al., 2024) leverages GPT (Radford, 2018; Radford et al., 2019) as its foundation due to its enhanced parallel processing capabilities and capacity. However, text-based LLMs such as GPT-3 (Brown, 2020) and Llama (Touvron et al., 2023) are essential for speech modeling. Once discrete audio representations are obtained, these large text models are trained to align with or enhance the original text embedding space, as explored in studies by (Zhang et al., 2023), (Fathullah et al., 2023), (Shu et al., 2023), and (Rubenstein et al., 2023). This trend of integrating textual LMs into speech modeling has become increasingly popular in recent research.
6 Limitations and Broader Impact
Limitations. In this work, we present the effectiveness of our proposed method, DM-Codec, based on the LibriSpeech dataset. Future research could investigate its performance across a variety of datasets and domains. Additionally, exploring the capabilities of DM-Codec in multilingual contexts would be valuable. Another limitation of our work is the absence of experiments with emerging LLMs. Currently, we focus solely on masked language models to derive representations. Further investigation into these decoder-based LLMs’ impact on DM-Codec can be studied and addressed.

### v2

@@ Appendix A Resources
Table 3: Significance Analysis on LibriSpeech Test Set. Significance analysis is conducted at $\alpha=0.05$ between LM and SM-guided DM-Codec (D), EnCodec (E), SpeechTokenizer (S), and FACodec (F). Comparisons are performed row vs. column (e.g., D vs. E, E vs. S). Results reveal DM-Codec consistently achieves significantly better scores in key metrics across all individual samples. A indicates significance, a denotes dominance, and a means no significance. Avg and Std mean the average and standard deviation of each score.
@@
Avg Std D E S F Avg Std D E S F Avg Std D E S F Avg Std D E S F
D 4.774 0.100 - 7.510 0.139 - 3.197 0.184 - 0.937 0.021 -
E 4.828 0.100 - 7.593 0.137 - 3.064 0.201 - 0.917 0.021 -
S 4.942 0.101 - 7.725 0.138 - 3.080 0.190 - 0.920 0.025 -
F 4.914 0.103 - 7.643 0.141 - 3.113 0.250 - 0.946 0.023 -
Appendix B Representation Alignment in LM Distillation
A core principle of our approach is that strict temporal alignment between text and acoustic representations is not necessary for effective contextual knowledge distillation. The discrete representations produced by the Residual Vector Quantizer (RVQ) inherently encode holistic information about the speech segment rather than temporally localized features. This characteristic enables the distillation of contextual representations into discrete tokens without reliance on strict temporal correspondence.
As illustrated in Figure 2, the RVQ outputs $\{Q_{1},\dots,Q_{T^{\prime}}\}$ capture comprehensive speech information across the entire utterance. Temporal alignment between these discrete representations and contextual representations (e.g., from BERT) or semantic representations (e.g., from HuBERT) is therefore not essential. The effectiveness of dimension-level semantic representation distillation without temporal alignment has been previously demonstrated in SpeechTokenizer (Zhang et al., 2024a), providing a robust theoretical foundation for this approach.
Moreover, vector quantizer output representations are not inherently aligned with time steps, but instead encode holistic speech information. This capability has been established in prior work (Zhang et al., 2024a; Huijben et al., 2024; Islam et al., 2024, 2023), which demonstrates the potential of vector quantization to discretize input data into intermediate representations that capture essential features across the feature dimension. Consequently, imposing a temporal alignment between vector quantizer outputs and hidden layer representations from language or semantic models would neither align with the methodological objectives nor enhance the efficacy of the proposed distillation approach.
To achieve effective contextual and semantic knowledge transfer, we employ a continuous distillation loss that maximizes cosine similarity at the feature dimension level between the selected RVQ layer outputs and the teacher representations across all time steps. Unlike conventional methods that rely on time-step-wise loss calculations (Zhang et al., 2024a; Chang et al., 2022), this dimension-level cosine similarity loss ensures that DM-Codec captures contextual and semantic knowledge through LM-Guided Distillation and Combined LM and SM-Guided Distillation mechanisms, without requiring strict temporal alignment.
In addition, we propose a [CLS]-token-based distillation strategy to address alignment. The [CLS] token encodes sequence-level holistic representations, capturing global contextual information from language models. Using these token representations, our method eliminates the need for fine-grained temporal alignment while preserving essential linguistic features. This complements the dimension-level distillation strategy by focusing on global sequence features, enabling the adaptability of our approach to scenarios where fine-grained alignment is infeasible or unnecessary.
As shown in Table 1, all DM-Codec variants, including DM-Codec (CLS), DM-Codec (LM), and DM-Codec (LM+SM) consistently outperform baseline models (EnCodec, SpeechTokenizer, and FACodec) in content preservation metrics (WER and WIL) while maintaining competitive performance in speech quality metrics (ViSQOL and STOI). These results corroborate the robustness of the proposed approach. By prioritizing the holistic integration of multimodal knowledge into discrete speech representations, DM-Codec achieves significant advancements in both content preservation and speech quality.
Appendix C Significance Analysis of Speech Tokenizer performance
We conducted a significance analysis at $\alpha=0.05$ on the LibriSpeech test-clean subset containing 2,620 samples. We follow the approach of Dror et al. (2019), to measure the stochastic dominance of DM-Codec over the baselines: EnCodec, SpeechTokenizer, and FACodec. Specifically, we computed inverse cumulative distribution functions (CDFs) for all reconstructed speech samples’ individual WER, WIL, ViSQOL, and STOI scores. Significance was evaluated using the $\epsilon$ value and categorized as: significantly better when $0.0<\epsilon\leq 0.5$, significantly dominant when $\epsilon=0.0$, and not significantly better when $\epsilon>0.5$. For this analysis, we selected DM-Codec (LM+SM), trained with combined LM and SM-guided distillation. To the best of our knowledge, we are the first to conduct significance analysis to measure the effectiveness of different speech tokenizers.
@@ Appendix C Significance Analysis of Speech Tokenizer performance
Results and Discussion: The results in Table 3 show that DM-Codec significantly outperforms the baselines in WER, WIL, ViSQOL, and STOI scores. The improved average values (4.774 WER, 7.510 WIL, 3.197 ViSQOL, 0.937 STOI) and consistent standard deviations (0.100 WER, 0.139 WIL, 0.184 ViSQOL, 0.021 STOI) further demonstrate the statistical significance. Notably, DM-Codec’s performance in WER and WIL underscores the importance of contextual representation distillation for enhanced speech reconstruction. Additionally, its strong performance in ViSQOL and STOI, especially over EnCodec, highlights the benefits of combining LM and SM distillation for retaining semantic-acoustic fidelity. While DM-Codec does not achieve significance over FACodec in terms of STOI, it significantly outperforms the baselines across all other metrics. Among the baseline, FACodec achieves significance over SpeechTokenizer, whereas EnCodec outperforms SpeechTokenizer in WER and WIL, SpeechTokenizer excels in ViSQOL and STOI over EnCodec.
@@
Figure 3: Effects of weights on combined distillation from Speech Model (SM) and Language Model (LM). Higher LM weight generally results in improved WER, suggesting its stronger contribution to content preservation. Here, $\lambda_{SM}$ is the weight added to SM and $\lambda_{LM}$ is the weight added to LM, where ($\lambda_{SM}$ + $\lambda_{LM}$ = 1).
Appendix D Ablation Studies
We conducted a comprehensive analysis of DM-Codec’s performance and the impact of each methodological choice in retaining contextual and semantic information through distillation. Unless otherwise stated, we use distillation for both LM and SM to the first Residual Vector Quantizer layer (RVQ-1) for comparison consistency and simplicity. The following ablation studies were conducted concurrently with a similar configuration and model design except for the explicitly noted changes.
D.1 Ablation Study: Impact of Combined Semantic Distillation
@@
$\mathcal{L}_{LS}=\frac{1}{2}\left(\lambda_{LM}\cdot\mathcal{L}_{L}+\lambda_{SM}\cdot\mathcal{L}_{S}\right)$ (13)
Results and Discussion: The experimental results are presented in Figure 3, showing the speech reconstruction results with WER scores for different weighted combinations. From the values, we notice a trend showing that incorporating LM representations generally improves WER, especially when LM distillation is dominant. The lowest WER score of 4.07 occurs with a weight of $\lambda_{LM}=0.8$ for LM, while $\lambda_{SM}=0.2$ for SM, highlighting the strong influence of LM distillation on capturing contextual information. A balanced weighting of $\lambda_{SM}=0.5$ and $\lambda_{LM}=0.5$ produces a WER of 4.18, confirming that distillation from both LM and SM is beneficial. However, as the weighting shifts more in favor of SM ($\lambda_{SM}>0.7$), WER deteriorates, reaching 4.83 when relying entirely on SM. This underscores that over-reliance on SM distillation compromises contextual accuracy in favor of raw speech features. Notably, the interaction between LM and SM weights plays a crucial role, as the combined distillation influences the overall WER beyond individual distillation contributions. For instance, the higher WER observed at $\lambda_{LM}=0.9$ compared to $\lambda_{LM}=0.3$ or $\lambda_{LM}=0.5$ highlights the importance of tuning both weights synergistically, rather than favoring one in isolation.
Table 4: Analysis of different RVQ layers effect on speech reconstruction. LM-guided distillation on RVQ-1 layer ensures greater content preservation, while SM-guided distillation on RVQ-1:8 layer is more effective at preserving semantic representation. LM-layer and SM-layer indicate the RVQ layer used for respective distillation. (LM) indicates LM-guided Distillation. (LM+SM) indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec (LM) RVQ-1 - 4.36 7.06 3.18 0.935
DM-Codec (LM) RVQ-1:8 - 4.23 6.94 3.12 0.929
DM-Codec (LM) RVQ-8 - 4.44 7.22 3.28 0.935
DM-Codec (LM+SM) RVQ-1 RVQ-1 4.18 6.84 3.13 0.933
DM-Codec (LM+SM) RVQ-1:8 RVQ-1 4.59 7.34 3.21 0.937
DM-Codec (LM+SM) RVQ-8 RVQ-1 4.49 7.24 3.30 0.938
DM-Codec (LM+SM) RVQ-1 RVQ-1:8 4.05 6.61 3.26 0.937
DM-Codec (LM+SM) RVQ-1 RVQ-8 4.39 7.08 3.33 0.939
D.2 Ablation Study: Impact of Distillation on Different RVQ Layers
@@
Table 5: Analysis of representation distillation from different models. BERT can be effectively combined with HuBERT or wav2vec 2.0, however, ELECTRA in LM-guided distillation outperforms BERT. (LM) indicates LM-guided Distillation. (LM+SM) indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec (LM) BERT - 4.36 7.06 3.18 0.935
DM-Codec (LM) ELECTRA - 4.12 6.63 3.10 0.936
DM-Codec (LM+SM) BERT HuBERT 4.18 6.84 3.13 0.933
DM-Codec (LM+SM) BERT wav2vec 2.0 4.13 6.77 3.15 0.942
DM-Codec (LM+SM) ELECTRA wav2vec 2.0 4.70 7.51 3.14 0.933
DM-Codec (LM+SM) ELECTRA HuBERT 4.67 7.58 2.94 0.932
D.3 Ablation Study: Impact of Different Models on Distillation
@@
Table 6: Analysis of different distillation layers representation on speech reconstruction. Average layer provides more comprehensive representations. (LM) indicates LM-guided Distillation. (LM+SM) indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
DM-Codec (LM) Average 4.36 7.06 3.18 0.935
DM-Codec (LM) Last 4.62 7.56 2.95 0.926
DM-Codec (LM) $9^{\text{th}}$ 4.75 7.80 2.88 0.925
DM-Codec (LM+SM) Average 4.18 6.84 3.13 0.933
DM-Codec (LM+SM) Last 4.68 7.55 3.03 0.933
DM-Codec (LM+SM) $9^{\text{th}}$ 4.52 7.43 3.00 0.933
Table 7: Analysis of different bit rates for speech reconstruction. DM-Codec (LM+SM) showcases its robustness at reduced bitrates, outperforming baselines at 3 kbps and maintaining competitive content preservation scores (WER, WIL) and superior speech quality (ViSQOL, STOI) at 1.5 kbps. $f_{s}$ represents the audio sample rate, and $f_{r}$ the codec frame rate. $\heartsuit$ means the results were reproduced using the official training code. $\diamondsuit$ means the results were obtained using official model checkpoints. (LM) indicates LM-guided Distillation method. (LM+SM) indicates combined LM and SM-guided Distillation method.
Model $f_{s}$ $f_{r}$ Bitrate WER $\downarrow$ WIL $\downarrow$ ViSQOL $\uparrow$ STOI $\uparrow$
DM-Codec (LM+SM) 16 kHz 50 Hz 3 kbps 4.29 7.04 3.070 0.928
DM-Codec (LM) 16 kHz 50 Hz 3 kbps 4.38 7.09 3.042 0.924
[…]
DM-Codec (LM) 16 kHz 50 Hz 1.5 kbps 6.14 10.13 2.644 0.880
DM-Codec (LM+SM) 16 kHz 50 Hz 1.5 kbps 6.19 10.16 2.662 0.894
[…]
DM-Codec (LM) 16 kHz 50 Hz 0.75 kbps 7.96 12.65 2.18 0.583
DM-Codec (LM+SM) 16 kHz 50 Hz 0.75 kbps 7.81 12.42 2.20 0.675
[…]
D.4 Ablation Study: Impact of Different Distillation Layer(s)
@@
D.5 Ablation Study: Impact of Low Bit Rate
We evaluated zero-shot speech reconstruction at reduced bitrates of 3kbps, 1.5kbps, and 0.75kbps. DM-Codec and the baseline SpeechTokenizer were trained on 16kHz sample rates, whereas the EnCodec baseline was trained on 24kHz sample rates. To achieve lower bitrates in DM-Codec and SpeechTokenizer, we limited RVQ levels: the first 6 RVQ layers for 3kbps, 3 RVQ layers for 1.5kbps, and 1 RVQ layer for 0.75kbps. For EnCodec, we kept the first 4 VQ layers for 3kbps and 2 VQ layers for 1.5kbps, and it does not support a 0.75kbps bitrate. In this ablation, DM-Codec variants used the first RVQ layer (RVQ-1) for LM distillation, while all RVQ layers (RVQ-1:8) were used for SM distillation. Table 7 presents the full results.
Results and Discussion: At a 3kbps bitrate, LM-guided DM-Codec (LM) maintains its performance and consistency, surpassing the baseline with scores of 4.38 WER, 7.09 WIL, 3.042 ViSQOL, and 0.924 STOI. Combined LM and SM-guided DM-Codec (LM+SM) further improves these scores to 4.29 WER, 7.04 WIL, 3.070 ViSQOL, and 0.928 STOI, outperforming all baselines.
At 1.5kbps and 0.75kbps, DM-Codec (LM) and DM-Codec (LM+SM) achieve scores comparable to SpeechTokenizer in WER and WIL but maintain superior speaker quality, achieving the best ViSQOL and STOI scores of 2.662 and 0.894 at 1.5kbps, and 2.20 and 0.675 at 0.75kbps, respectively. We hypothesize that the slight performance degradation in WER and WIL at lower bitrates is due to the loss of contextual representation, as the reduced bandwidth limits the model’s ability to fully utilize nuanced contextual details incorporated into DM-Codec through distillation.
@@ D.5 Ablation Study: Impact of Low Bit Rate
Table 8: Analysis of different distillation axes for speech reconstruction. Distillation at the feature dimension is more robust and enhances reconstruction. D-axis indicates the proposed distillation along the feature dimension. T-axis indicates distillation along the time axis. (LM) indicates LM-guided Distillation. (LM+SM) indicates combined LM and SM-guided Distillation. Bold highlights the best result and underline the second-best result.
Model Axis WER $\downarrow$ WIL $\downarrow$ ViSQOL $\uparrow$ STOI $\uparrow$
DM-Codec (LM) D-axis 4.36 7.06 3.18 0.935
DM-Codec (LM) T-axis 4.76 7.82 2.92 0.925
DM-Codec (LM+SM) D-axis 4.05 6.61 3.26 0.937
DM-Codec (LM+SM) T-axis 4.76 7.76 2.86 0.926
Table 9: Analysis of different distillation targets for speech reconstruction. (LM) indicates LM-guided distillation with averaged token representations, (CLS) indicates [CLS]-token-based distillation, (Baseline) refers to DM-Codec without any distillation, and (Word Embedding) indicates static BERT word embeddings. Contextual LM representations (LM, CLS) consistently outperform both the Baseline and static Word Embeddings. Bold highlights the best result and underline the second-best result.
Model WER $\downarrow$ WIL $\downarrow$ ViSQOL $\uparrow$ STOI $\uparrow$
DM-Codec (Baseline) 4.97 8.02 2.95 0.935
DM-Codec (LM) 4.36 7.06 3.18 0.935
DM-Codec (CLS) 4.47 7.08 3.12 0.926
DM-Codec (Word Embedding) 5.27 8.58 2.89 0.931
D.6 Ablation Study: Impact of Distillation Axis
We experiment with different axes for feature distillation. In Eqn. 1 and 2, we calculate the distillation loss by maximizing cosine similarity along the feature dimension axis (D-axis). To evaluate the impact of an alternative approach, we conduct this ablation study by maximizing cosine similarity along the time dimension axis (T-axis). For this ablation, we modify Eqn. 1 and 2 by computing cosine similarity across each time step instead of each feature dimension. Table 8 presents the complete results. We select D-axis DM-Codec variants where the first Residual Vector Quantizer layer (RVQ-1) was used for LM distillation and all RVQ layers (RVQ-1:8) for SM distillation and retrain with the T-axis formulation.
Results and Discussion: In LM-guided distillation, the D-axis consistently achieves better performance across all metrics, with scores of 4.36 WER, 7.06 WIL, 3.18 ViSQOL, and 0.935 STOI, outperforming the relatively lower scores of the T-axis. Similarly, for combined LM and SM-guided distillation, the D-axis achieves superior results with 4.05 WER, 6.61 WIL, 3.26 ViSQOL, and 0.937 STOI compared to the T-axis. These results demonstrate that calculating cosine similarity along the feature dimension axis (D-axis) is empirically more robust and effective. In contrast, cosine similarity along the time axis (T-axis) fails to capture contextual and semantic information. The identical WER of 4.76 for both LM and LM+SM distillation on the T-axis further highlights its limitations. However, our D-axis approach effectively incorporates richer information within each feature dimension, enhancing the model’s capacity to capture nuanced patterns. This leads to improved speech reconstruction, validating our proposed method’s robustness.
D.7 Ablation Study: Impact of Word Embedding Distillation Targets
We conducted an experiment using static BERT word embeddings as the distillation target $L_{LM}$, referred to as DM-Codec (Word Embedding). Table 9 compares different distillation targets.
Results and Discussion: Both DM-Codec (LM) and DM-Codec (CLS), which distill contextual token representations from the LM, clearly outperform DM-Codec (Baseline) and DM-Codec (Word Embedding). In particular, DM-Codec (LM) achieves the best overall balance, with strong WER (4.36), WIL (7.06), ViSQOL (3.18), and STOI (0.935).
By contrast, DM-Codec (Word Embedding), which relies on static embeddings, performs the worst across intelligibility metrics (WER 5.27, WIL 8.58) and shows limited improvement in perceptual quality (ViSQOL 2.89, STOI 0.931). This suggests that static word embeddings, lacking contextual representation, are insufficient as distillation targets for high-quality speech reconstruction. Overall, these results underscore the importance of contextualized LM representations in guiding effective distillation.
Appendix E DM-Codec Model Components
Residual Vector Quantizers. The Residual Vector Quantizer (RVQ) plays a central role in our tokenization process, quantizing the encoder’s outputs. Our implementation is inspired by the training procedures described in Encodec (Défossez et al., 2022) and SpeechTokenizer (Zhang et al., 2024a). The RVQ projects input vectors to the most similar entry in a codebook, and the residual is calculated and processed in subsequent quantization steps, each utilizing a different codebook. The codebook entries are updated using an exponential moving average (EMA) with a decay rate of 0.99 for the matched item, while unmatched entries are replaced by candidates from the current batch. To ensure proper gradient flow during training, we employ a straight-through estimator. A commitment loss is also computed and added to the total training loss to promote stability.
Appendix F Human Evaluation Methodology
To evaluate the quality and effectiveness of our approach, we conducted human evaluations using Mean Opinion Score (MOS) and Similarity Mean Opinion Score (SMOS) metrics, following methodologies established in prior works such as SpeechTokenizer and Vall-E. The study was conducted under an approved Institutional Review Board (IRB) protocol to ensure ethical compliance and participant safety. A total of 50 proficient English speakers, comprising graduate and undergraduate students, were selected as evaluators based on their high language comprehensibility. These participants volunteered for the evaluation, were briefed on the study’s purpose, and were provided no information that could bias their judgments.
The evaluation process involved each participant rating batches of fully anonymized and randomized speech samples via a web-based survey interface, with clear and standardized guidelines to ensure consistent and unbiased scoring. Each batch contained 16 samples, including outputs from both our proposed models and baseline systems. For the speech reconstruction task, participants rated the perceptual quality of the speech samples using the MOS, based on criteria such as naturalness, intelligibility, and clarity, employing a 5-point Likert scale, where higher scores indicated superior quality. For the text-to-speech evaluation, participants provided two distinct ratings: MOS, to measure the overall naturalness of the generated speech, and SMOS, to evaluate the similarity of the generated speech to the original speaker’s voice, both rated on a 1-to-5 scale with 1-point increments. To enhance reliability and mitigate individual evaluator bias, each sample was rated by multiple participants. Figure 5 shows the interface view and instructions given to participants.
Interface View 1
Interface View 2
Interface View 3
Interface View 4
Interface View 5
Figure 5: Web-based survey interface and questionnaire used for human evaluation of samples.
Appendix G Broader Impact and Potential Risks